The length of a rectangle is twice as its width. Its area is 6 more than the width. Find the dimension of the rectangle.
Answer by ankor@dixie-net.com(15658)

The length of a rectangle is twice as its width.
L = 2W
:
Its area is 6 more than the width. Find the dimension of the rectangle.
L*W = W+6
:
Substitute 2W for L in the above equation, Find W:
2W * W = W + 6
2W^2 - W - 6 = 0; a quadratic equation
:
Factos to:
(2W + 3)(W - 2) = 0
2W = -3
W = -1.5
and
W = +2 is the width
:
That makes the Length = 4
:
Check:
Area = Width + 6
2*4 = 2 + 6